# Approximation of multipartite quantum states: revised version with new applications

### v1

@@ 2.2 Simple sufficient condition for the FA-property
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}g_{i}<+\infty\quad\textrm{and}\quad\lim_{\beta\to 0^{+}}\left[\sum_{i=1}^{+\infty}e^{-\beta g_{i}}\right]^{\beta}=1.$ (12)
@@
Proof. Let $\rho=\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}|\varphi_{i}\rangle\langle\varphi_{i}|$ be a spectral decomposition of $\rho$.
If (13) holds then there exists an increasing sequence $\{c_{i}\}_{i=1}^{+\infty}$ of positive numbers tending to +$\infty$+ ∞ such that
@@
It is easy to see that the operator
@@
satisfies condition (8). By Lemma 2.1 the operator $H^{2}$ satisfies condition (10). Since (14) implies that $\mathrm{Tr}H^{2}\rho<+\infty$, the state $\rho$ has the FA-property by the Theorem in [32, Section 2]. $\Box$
@@
Indeed, it is easy to see that any state with the spectrum $\{c[i\ln i\ln^{p}(\ln i)]^{-1}\}_{i=10}^{+\infty}$, $p>1$, where $c$ is a normalizing constant, satisfies condition (13) but does not satisfy condition (15).
@@
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}\lambda^{\rho}_{i}<+\infty\;\Rightarrow\;\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}i<+\infty\;\Rightarrow\;\rho\in\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H})\;\Rightarrow\;S(\rho)<+\infty,$ (16)
where $\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H})$ denotes the subset of states having the FA-property (the first implication in (16) follows from the inequality $\lambda^{\rho}_{i}\leq 1/i,\;\forall i$, the third one is mentioned in the Theorem in [32, Section 2]). In [2] it is shown that the third implication in (16) cannot be replaced by """"$\Leftrightarrow$"" ⇔ " by constructing examples of states with finite entropy which do not have the FA-property. We may conjecture (until proven otherwise) that the second implication in (16) can be replaced by """"$\Leftrightarrow$"" ⇔ ", i.e. that (13) is a criterion of the FA-property of a state with the spectrum $\{\lambda^{\rho}_{i}\}$ arranged in the non-increasing order.
@@ 2.3 On uniform continuity under the energy-type constraints
Assume that $H_{A_{1}}$,…,$H_{A_{m}}$ are arbitrary positive densely defined operators on the spaces $\mathcal{H}_{A_{1}}$,…,$\mathcal{H}_{A_{m}}$, $m\leq n$. For a given function $f$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, $\varepsilon\in(0,1]$ and $E>0$ introduce the quantity
@@
$\lim_{\varepsilon\to 0}\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})=0\quad\forall E>E_{0}$
for arbitrary positive operators $H_{A_{1}}$,…,$H_{A_{m}}$ satisfying condition (8) (resp. condition (10)), where $mE_{0}$ is the sum of the minimal eigenvalues of the operators $H_{A_{1}}$,…,$H_{A_{m}}$.
Let $\mathcal{F}$ be a family of functions from the class $I_{n}^{m}$ (resp. $II_{n}^{m}$). We will say that $\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})$ is a common faithful $I_{n}^{m}$-type (resp. $II_{n}^{m}$-type) continuity bound for the family $\mathcal{F}$ provided that
$B_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})\leq\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})\quad\forall f\in\mathcal{F},\;E>E_{0},\;\varepsilon\in(0,1]$
and
$\lim_{\varepsilon\to 0}\widehat{B}_{\mathcal{F}}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})=0\quad\forall E>E_{0}$
for arbitrary positive operators $H_{A_{1}}$,…,$H_{A_{m}}$ satisfying condition (8) (resp. condition (10)).
Among characteristics of a composite infinite-dimensional quantum system $A_{1}...A_{n}$ there are many that belong to the classes $I_{n}^{m}$ and $II_{n}^{m}$ for some $m\leq n$ [^7] It follows from the definitions that if $f\in I_{n}^{m}$ (resp. $f\in II_{n}^{m}$) for some $m<n$ then $f\in I_{n}^{u}$ (resp. $f\in II_{n}^{u}$) for any $u=\overline{m,n}$. For example, the von Neumann entropy of a state of $A_{1}$ and the quantum conditional entropy $S(A_{1}|A_{2})$ belong, respectively, to the classes $I_{1}^{1}$ and $I^{1}_{2}$. The $I_{1}^{1}$-type and $I^{1}_{2}$-type continuity bounds for these characteristics are obtained by Winter in [41]. The $I^{1}_{2}$-type continuity bounds for the quantum mutual information $I(A_{1}\!:\!A_{2})$, for the entanglement of formation $E_{F}$ and for the squashed entanglement $E_{sq}$ in a bipartite system $A_{1}A_{2}$ obtained in [27] by using Winter’s technique show that these characteristics belong to the class $I^{1}_{2}$. Moreover, the last two entanglement measures have a common faithful $I^{1}_{2}$-type continuity bound. The regularized entanglement of formation $E^{\infty}_{F}$ also belongs to the class $I^{1}_{2}$: the $I^{1}_{2}$-type continuity bounds for $E^{\infty}_{F}$ obtained by using Winter’s technique is presented in Proposition 18 in [33]. [^8] We do not distinguish between the discrete and continuous versions of $E_{F}$ and $E^{\infty}_{F}$, since these versions coincide at any state $\omega\in\mathfrak{S}(\mathcal{H}_{A_{1}A_{2}})$ such that at least one of the states $\omega_{A_{1}}$ and $\omega_{A_{2}}$ has finite energy w.r.t. a Hamiltonian satisfying condition (8) [33, Section 4.4.2].
@@
where $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$, $s=1,2,3$, are the classes of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ introduced in [31, 33] which contain many important characteristics of composite quantum systems. [^9] A non-complete list of these characteristics and detailed description of the classes $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$, $s=1,2,3$, can be found in the Appendix. Moreover, Theorem 6 in [33] implies that all functions in the family
$\mathcal{F}(C,D)=\widehat{L}^{m}_{n}(C,D)\cup N^{m}_{n,1}(C,D)\cup N^{m}_{n,2}(C,D)\cup N^{m}_{n,3}(C,D),\quad C,D\geq 0,$
@@
$\!\widehat{B}_{\mathcal{F}(C,D)}(\varepsilon,E|H_{A_{1}},..,H_{A_{m}})=C\sqrt{\varepsilon(2-\varepsilon)}F_{H_{A_{1}},..,H_{A_{m}}}\!\!\left[\frac{2mE}{\varepsilon(2-\varepsilon)}\right]+Dg(\sqrt{\varepsilon(2-\varepsilon)}),$ (19)
@@
$F_{H_{A_{1}},...,H_{A_{m}}}(E)=\sup\left\{S(\rho)\,\left|\,\rho\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{m}}),\,\sum_{k=1}^{m}\mathrm{Tr}H_{A_{k}}\rho_{A_{k}}\leq E\right.\right\}$ (20)
@@
$g(x)=(x+1)h_{2}\!\left(\frac{x}{x+1}\right)=(x+1)\ln(x+1)-x\ln x,\;\,x>0,\quad g(0)=0.$ (21)
The faithfulness of the common continuity bound (19) for the family $\mathcal{F}(C,D)\subset II_{n}^{m}$ follows from the fact that the validity of condition (10) for all the operators $H_{A_{1}},...,H_{A_{m}}$ implies that
@@
We will see in the next Section 3 that belonging of a given characteristic to the class $II_{n}^{m}$ allows us to realise effective finite-dimensional approximation of this characteristic. If, moreover, this characteristic belongs to the class $I_{n}^{m}$ then we may relax some requirements in such approximation (briefly speaking, to replace the FA-property of marginal states by the finiteness of the von Neumann entropy of these states).
Practically, to show that a given characteristic $f$ belongs to the class $II_{n}^{m}$ we have either to use general relation (18) or to construct an explicit continuity bound $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for $f$ which tends to zero as $\varepsilon\to 0$ for arbitrary positive operators $H_{A_{1}}$,….,$H_{A_{m}}$ satisfying condition (10). More efforts are required to show that a given characteristic $f$ belongs to the class $I_{n}^{m}$: it is necessary to construct an explicit continuity bound $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for $f$ which tends to zero as $\varepsilon\to 0$ for arbitrary positive operators $H_{A_{1}}$,….,$H_{A_{m}}$ satisfying condition (8). Sometimes one can use the following simple
@@
Example 1. The one-way classical correlation $C_{B}$ and the quantum discord $D_{B}$ in a bipartite infinite-dimensional system $AB$ with measured system $B$ (cf.[7, 20, 35]) belong to class $II_{2}^{1}$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$. This follows from (18), since $C_{B}\in\widehat{L}_{2}^{1}(1,2)$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$ and $D_{B}\in\widehat{L}_{2}^{1}(C,2)$, where $C=2$ in the setting $A_{1}=A,A_{2}=B$ and $C=1$ in the setting $A_{1}=B,A_{2}=A$ [33, Section 4.3]. The $II_{2}^{1}$-type continuity bounds for the regularized one-way classical correlation $C^{\infty}_{B}$ presented in [33, Proposition 16] implies that $C^{\infty}_{B}\in II_{2}^{1}$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$.
More detailed analysis shows that the functions $C_{B}$, $C^{\infty}_{B}$ and $D_{B}$ belong to class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$. Indeed, by the Koashi-Winter relations (cf.[13]) we have
@@
By applying the faithful $I_{1}^{1}$-type continuity bound for the entropy and the $I_{2}^{1}$-type continuity bounds for the entanglement of formation and its regularization mentioned before (18) it is easy to show that the functions $\omega\mapsto S(\omega_{A})$, $\omega\mapsto E_{F}(\omega_{AC})$ and $\omega\mapsto E^{\infty}_{F}(\omega_{AC})$ on $\mathfrak{S}(\mathcal{H}_{ABC})$ have faithful $I_{3}^{1}$-type continuity bounds in the setting $A_{1}=A,A_{2}=B,A_{3}=C$.
Since for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{AB})$ such that $\frac{1}{2}\|\rho-\sigma\|_{1}\leq\varepsilon$ there exist pure states $\hat{\rho}$ and $\hat{\sigma}$ in $\mathfrak{S}(\mathcal{H}_{ABC})$ such that $\frac{1}{2}\|\hat{\rho}-\hat{\sigma}\|_{1}\leq\sqrt{\varepsilon(2-\varepsilon)}$ [9, 39], using the continuity bounds mentined before, the expressions in (22) and Lemma 2.3 one can prove that $C_{B}$ and $C^{\infty}_{B}$ belong to class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$.
To prove that $D_{B}$ belongs to class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$, it suffice to note that $D_{B}=I(A\!:\!B)-C_{B}$ and that $I(A\!:\!B)$ belongs to class $I_{2}^{1}$ in the same setting (as mentioned before) and to use Lemma 2.3.
The question of whether the functions $C_{B}$, $C^{\infty}_{B}$ and $D_{B}$ belong to class $I_{2}^{1}$ in the setting $A_{1}=B,A_{2}=A$ remains open. $\Box$
@@
$I(A_{1}\!:\!...\!:\!A_{n})_{\omega}=I(A_{n-1}\!:\!A_{n})_{\omega}+I(A_{n-2}\!:\!A_{n-1}A_{n})_{\omega}+...+I(A_{1}\!:\!A_{2}...A_{n})_{\omega}$ (23)
valid for any state $\omega\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ and Lemma 2.3 it is easy to show that the $n$-partite quantum mutual information $I(A_{1}\!:\!...\!:\!A_{n})$ belongs to class $I_{n}^{n-1}$.
By the same way one can show that the (extended) quantum conditional mutual information $I(A_{1}\!:\!...\!:\!A_{n}|A_{n+1})$ (defined by the rule described in [33, Section 4.2]) belongs to class $I_{n+1}^{n-1}$ (since $I(A_{1}\!:\!A_{2}|A_{3})$ lies in $I_{3}^{1}$ due to the $I_{3}^{1}$-type continuity bound for this characteristic obtained in [27]). $\Box$

### v2

@@ 2.2 Simple sufficient condition for the FA-property
Proof. W.l.o.g. we may assume that $\rho$ is an infinite rank state.
Let $\rho=\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}|\varphi_{i}\rangle\langle\varphi_{i}|$ be a spectral decomposition of $\rho$. If (19) holds then there exists an increasing sequence $\{c_{i}\}_{i=1}^{+\infty}$ of positive numbers tending to +$\infty$+ ∞ such that
@@
It is easy to see that the positive operator
@@
on $\mathcal{D}(H)=\{\psi\in\mathrm{supp}\rho\,|\,\sum_{i=1}^{+\infty}c_{i}\ln^{2}i\hskip 1.0pt|\langle\varphi_{i}|\psi\rangle|^{2}<+\infty\}$ satisfies condition (15). By Lemma 2.1 the operator $H^{2}$ satisfies condition (17). Since (20) implies that $\mathrm{Tr}H^{2}\rho<+\infty$, the state $\rho$ has the FA-property by the Theorem in [4, Section 2]. $\Box$
@@
Indeed, it is easy to see that any state with the spectrum $\{c[i\ln^{p}(\ln i)\ln^{3}i\hskip 1.0pt]^{-1}\}_{i=10}^{+\infty}$, $p>1$, where $c$ is a normalizing constant, satisfies condition (19) but does not satisfy condition (21).
@@
$\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}\lambda^{\rho}_{i}<+\infty\;\Rightarrow\;\sum_{i=1}^{+\infty}\lambda^{\rho}_{i}\ln^{2}i<+\infty\;\Rightarrow\;\rho\in\mathfrak{S}_{\rm\textsf{FA}\!}(\mathcal{H})\;\Rightarrow\;S(\rho)<+\infty.$ (22)
The first implication in (22) follows from the inequality $\lambda^{\rho}_{i}\leq 1/i,\;\forall i$, the third one is mentioned in the Theorem in [4, Section 2]. In [5] it is shown that the third implication in (22) cannot be replaced by """"$\Leftrightarrow$"" ⇔ ". S.Becker proved recently that the second implication in (22) cannot be replaced by """"$\Leftrightarrow$"" ⇔ " as well [22]. Therefore, the question of finding a simple criterion for the FA-property remains open.
@@ 2.3 On uniform continuity under the energy-type constraints
Assume that $H_{A_{1}}$,…,$H_{A_{m}}$ are arbitrary positive operators on the spaces $\mathcal{H}_{A_{1}}$,…,$\mathcal{H}_{A_{m}}$, $m\leq n$. For a given function $f$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ taking a finite value at any state $\rho$ with finite ”energy” $\sum_{k=1}^{m}\mathrm{Tr}H_{A_{k}}\rho_{A_{k}}$, given $\varepsilon\in(0,1]$ and $E>0$ introduce the quantity
@@
$\lim_{\varepsilon\to 0^{+}}\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},...,H_{A_{m}})=0\quad\forall E>E_{0}/m$
for arbitrary positive operators $H_{A_{1}}$,…,$H_{A_{m}}$ satisfying condition (15) (resp. condition (17)), where $E_{0}$ is the sum of the minimal eigenvalues of the operators $H_{A_{1}}$,…,$H_{A_{m}}$.
[…]
@@
where $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$, $s=1,2,3$, are the classes of functions on $\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$ introduced in [15, 26] which contain many important characteristics of composite quantum systems. Moreover, Theorem 6 in [26] implies that all functions in
$\widehat{L}^{m}_{n}(C,D)\cup N^{m}_{n,1}(C,D)\cup N^{m}_{n,2}(C,D)\cup N^{m}_{n,3}(C,D),\quad C,D\geq 0,$
@@
$\!\widehat{B}_{C,D}(\varepsilon,E|H_{A_{1}},..,H_{A_{m}})=C\sqrt{\varepsilon(2-\varepsilon)}F_{H_{A_{1}},..,H_{A_{m}}}\!\!\left[\frac{2mE}{\varepsilon(2-\varepsilon)}\right]+Dg(\sqrt{\varepsilon(2-\varepsilon)}),$ (25)
@@
$F_{H_{A_{1}},...,H_{A_{m}}}(E)=\sup\left\{S(\varrho)\,\left|\,\varrho\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{m}}),\,\sum_{k=1}^{m}\mathrm{Tr}H_{A_{k}}\varrho_{A_{k}}\leq E\right.\right\}$
@@
$g(x)\doteq(x+1)h_{2}\!\left(\frac{x}{x+1}\right)=(x+1)\ln(x+1)-x\ln x,\;\,x>0,\quad g(0)=0.$
The faithfulness of the common continuity bound (25) follows from the fact that the validity of condition (17) for all the operators $H_{A_{1}},...,H_{A_{m}}$ implies that
@@
We will see in the next Section 3 that for any given characteristic $f$ from the class $II_{n}^{m}$ one can realise its effective finite-dimensional approximation which allows us to prove nontrivial properties of $f$. If, moreover, this characteristic belongs to the class $I_{n}^{m}$ then we may relax some requirements in such approximation (briefly speaking, to replace the FA-property of marginal states by the finiteness of the von Neumann entropy of these states).
Practically, to show that a given characteristic $f$ belongs to the class $II_{n}^{m}$ we have either to use general relation (24) or to directly prove the existence of a continuity bound $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for $f$ which tends to zero as $\varepsilon\to 0$ for arbitrary positive operators $H_{A_{1}}$,….,$H_{A_{m}}$ satisfying condition (17). More efforts are required to show that a given characteristic $f$ belongs to the class $I_{n}^{m}$: it is necessary to prove the existence of a continuity bound $\widehat{B}_{f}(\varepsilon,E|H_{A_{1}},....,H_{A_{m}})$ for $f$ which tends to zero as $\varepsilon\to 0$ for arbitrary positive operators $H_{A_{1}}$,….,$H_{A_{m}}$ satisfying condition (15). Sometimes one can use the following simple
@@
Example 1. The one-way classical correlation $C_{B}$ and the quantum discord $D_{B}$ in a bipartite infinite-dimensional system $AB$ with measured system $B$ (cf.[28, 29, 30]) belong to the class $II_{2}^{1}$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$. This follows from (24), since $C_{B}\in\widehat{L}_{2}^{1}(1,2)$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$ and $D_{B}\in\widehat{L}_{2}^{1}(C,2)$, where $C=2$ in the setting $A_{1}=A,A_{2}=B$ and $C=1$ in the setting $A_{1}=B,A_{2}=A$ [26, Section 4.3]. The $II_{2}^{1}$-type continuity bounds for the regularized one-way classical correlation $C^{\infty}_{B}$ presented in [26, Proposition 16] implies that $C^{\infty}_{B}\in II_{2}^{1}$ in both settings $A_{1}=A,A_{2}=B$ and $A_{1}=B,A_{2}=A$.
A more detailed analysis shows that the functions $C_{B}$, $C^{\infty}_{B}$ and $D_{B}$ belong to the class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$. Indeed, by the Koashi-Winter relations (cf.[31]) we have
@@
By applying the faithful $I_{1}^{1}$-type continuity bound for the entropy and the faithful $I_{2}^{1}$-type continuity bounds for the entanglement of formation and its regularization mentioned before (24) it is easy to show that the functions $\,\omega\mapsto S(\omega_{A})$, $\omega\mapsto E_{F}(\omega_{AC})$ and $\,\omega\mapsto E^{\infty}_{F}(\omega_{AC})$ on $\mathfrak{S}(\mathcal{H}_{ABC})$ have faithful $I_{3}^{1}$-type continuity bounds in the setting $A_{1}=A,A_{2}=B,A_{3}=C$ (because the trace norm does not increase under partial trace).
Since for any states $\rho$ and $\sigma$ in $\mathfrak{S}(\mathcal{H}_{AB})$ such that $\frac{1}{2}\|\rho-\sigma\|_{1}<\varepsilon$ there exist pure states $\hat{\rho}$ and $\hat{\sigma}$ in $\mathfrak{S}(\mathcal{H}_{ABC})$ such that $\frac{1}{2}\|\hat{\rho}-\hat{\sigma}\|_{1}\leq\sqrt{2\varepsilon}$ [2, 9], using the continuity bounds mentined before, the expressions in (26) and Lemma 2.3 one can prove that $C_{B}$ and $C^{\infty}_{B}$ belong to the class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$.
To prove that $D_{B}$ belongs to the class $I_{2}^{1}$ in the setting $A_{1}=A,A_{2}=B$, it suffices to note that $D_{B}=I(A\!:\!B)-C_{B}$ and that $I(A\!:\!B)$ belongs to the class $I_{2}^{1}$ in the same setting (as mentioned before (24)) and to use Lemma 2.3.
The question of whether the functions $C_{B}$, $C^{\infty}_{B}$ and $D_{B}$ belong to the class $I_{2}^{1}$ in the setting $A_{1}=B,A_{2}=A$ remains open. $\Box$
@@
$I(A_{1}\!:\!...\!:\!A_{n})_{\omega}=I(A_{n-1}\!:\!A_{n})_{\omega}+I(A_{n-2}\!:\!A_{n-1}A_{n})_{\omega}+...+I(A_{1}\!:\!A_{2}...A_{n})_{\omega}$
(valid for any state $\omega\in\mathfrak{S}(\mathcal{H}_{A_{1}..A_{n}})$) and Lemma 2.3 it is easy to show that the $n$-partite quantum mutual information $I(A_{1}\!:\!...\!:\!A_{n})$ belongs to the class $I_{n}^{n-1}$.
By the same way one can show that the (extended) quantum conditional mutual information $I(A_{1}\!:\!...\!:\!A_{n}|A_{n+1})$ (defined by the rule described in [26, Section 4.2]) belongs to the class $I_{n+1}^{n-1}$ (since $I(A_{1}\!:\!A_{2}|A_{3})$ lies in $I_{3}^{1}$ due to the $I_{3}^{1}$-type faithful continuity bound for this characteristic obtained in [24]). $\Box$